a rocket is launched from an 18-foot cliff with an initial velocity of 80 feet per second. the height above the ground of the rocket t seconds after launch is given by the equation h= -16^2+80t+18. hjow long after the rocket is launched will it hit the ground? round your answer to the nearest tenth.
Answer by Alan3354(69443)

a rocket is launched from an 18-foot cliff with an initial velocity of 80 feet per second. the height above the ground of the rocket t seconds after launch is given by the equation h= -16^2+80t+18. hjow long after the rocket is launched will it hit the ground? round your answer to the nearest tenth.
-------------
It's when h = 0.
h= -16^2+80t+18 = 0
Solve for t.
---------------
Why is this called a rocket? It has no engine, doesn't accelerate.